# Sensitivity and control of a 6-axis fused-silica seismometer

### v1

@@ Abstract
We present a pair of seismometers capable of measurement in all six axes of rigid motion. The vacuum-compatible devices implement compact interferometric displacement sensors to surpass the sensitivity of typical electrical readout schemes. Together with the capability to subtract the sensitivity-limiting coupling of ground tilt into horizontal motion, our seismometers can widen the sensing band towards mHz frequencies. This has notable applications across a range of fields requiring access to low-frequency signals, such as seismology and climate research. We particularly highlight their potential application in gravitational-wave observatories (LIGO) for observation of intermediate-mass black holes (). The sensors are based on a near-monolithic fused-silica design consisting of a fused-silica mass and fibre, showing improved stability and robustness to tilt drifts, alignment, and control compared to all-metal or mixed metal-silica designs. We demonstrate tilt sensitivity that surpasses the best commercial alternatives in a significantly reduced footprint compared to our previous iterations of these sensors.
@@ I Introduction
However, even more precise sensors are required for fundamental physics research, including observation of gravitational waves. In this case inertial sensors are used within a broader feedback scheme to sense and control the motion of platforms and larger devices—the technique of inertial isolation. Inertial isolation can also be seen as an essential enabling technology for quantum sensing and computing develeopment, and applications in industrial manufacturing processes.
@@
In this paper, we present the design and performance of two six-axis inertial sensors based on ideas laid out in Ref. [12]. The purpose of these devices is to address the above limitations through (i) taking advantage of compact interferometric sensors with sub-picometre displacement sensitivity, (ii) creating an unconstrained system with full six-axis freedom of motion, and (iii) achieving low resonant frequencies along key DoFs. The ultimate goal is to provide better sensitivity across the full sensing band and to extend this band towards lower frequencies. The final designs and performance presented here follow extensive prior development of numerous metal precursors [13, 14, 15]. We enhanced the sensitivity and robustness of the seismometers by (i) utilising fused silica instead of aluminium to suppress drifts, (ii) making the devices compact to fit into LIGO-type inertial isolation platforms, (iii) developing compact interferometers with deep-frequency modulation readout, and (iv) introducing magnetoguns to correct the angular position of the mass.
@@
These devices are versatile and thus can benefit a multitude of disparate fields of research. The sensor can be used directly as a seismometer, where it can detect low-frequency signals and thus obtain more complete information about seismic transients, as well as to provide insight into low-frequency seismic phenomena. The importance of low frequency seismic information has been demonstrated in earthquake analysis [16], hydrocarbon surveys [17], and climate research [18]. Better sensitivity can also reduce the amplitude of acoustic excitations generated during active seismic surveys, which would be less disruptive to local wildlife [19].
@@
Our inertial sensors can be integrated into actively controlled platforms and provide a vibration-free environment for general use. Most prominently for us is its use in gravitational wave (GW) detectors [20, 21, 22, 23] (see Section II). Additionally, our sensor can be applied to suppress vibrations of cryogenic systems [24, 25], electron microscopes [26] and used as a test-bed for space-borne instruments [27]. Outside of research applications, the sensors can be used in ultra-precision manufacturing [28] and photolithography [29, 30].
@@
Finally, the sensors we developed can be used in the role of a torsion balance. Torsion balances have a strong history in gravity measurements with recent examples including Bosonic dark matter searches [31], and Newtonian noise measurement [32]. We have recently adapted a precursor sensor to investigate semi-classical gravity models [33, 34, 35], with results currently in preparation. It is also possible to apply such a device in air-borne gravity gradiometry [36, 37] and short-range force measurements investigating the Casimir effect in meta-materials [38, 39, 40].
@@ II LIGO Application
The Advanced LIGO [41] and Advanced Virgo [42] detectors have successfully yield dozens of GW sources [43, 44, 45] since the first GW transient detection in 2015 [46]. However, the duty cycle of the detectors, as well as their sensitivity below 25 Hz [47, 48], is significantly limited by ground vibrations. The reduced duty cycle leads to the loss of otherwise detectable sources and the narrower detection band [49] limits the accumulation of signal-to-noise ratio. This degrades the parameter estimation and sky localisation, and limits the early warning time [50] for follow-up and multi-messenger observations [51]. Widening the band also increases the prospects to detect and study intermediate-mass black holes [52].
@@
Commercial sensors are limited by the noise of their inductive/capacitive readout exacerbated by the degrading response below . Horizontal single-axis sensors further suffer from tilt-horizontal cross-couplings. Ground tilt around the perpendicular horizontal axis couples into the readout with a factor of [60], where is the angular frequency, which masks the true horizontal signal at low frequencies. One solution is to make an independent estimate of the ground tilt with two spatially separated vertical sensors and subtract it from the horizontal sensor. This method is limited by the relatively short, metre-scale coherence length of the ground tilt and the space constraints on the typical isolated platform. It also relies on multiple stages of coherent subtraction, up to a common-mode rejection ratio of 1000, which is highly susceptible to calibration errors and systematics.
A wide range of custom inertial sensors have been demonstrated or proposed in the literature. Sensors with minimal tilt coupling have been proposed [61] to bypass the cross-coupling effect and dedicated rotation sensors [62, 63, 64] have been developed to independently isolate the platforms from ground tilts. Improved sensors based around optical readout have also been proposed [65, 66, 67, 68, 69, 70]. Our focus has been on the development of a single six-axis seismometer that overcomes all of the key limitations simultaneously in a single device.
@@ III Mechanical Design
An essential component of our design is the single fused-silica suspension fibre—the same fibre in both tests. This fibre was pulled at the University of Glasgow similarly to the fibres on our precursor designs, using the same technique developed for the Advanced LIGO quadruple suspensions [72]. We observe the beneficial linear, elastic behaviour of the fibre, which exhibits no detectable hysteresis as witnessed in the RZ torsional motion. This, among other properties of fused silica, is a key motivation for our move towards a monolithic fused-silica design, replacing the previous aluminium test mass.
The fused-silica test mass is wheel-shaped with most of the mass concentrated at the edges of the ring to maximise the moment of inertia for the given mass. This allows us to reduce the actuation strength needed to control the linear DoFs and use a thinner fibre. The higher moment of inertia reduces the tilt resonances and thus leads to better tilt sensitivity. Here, we see the benefits of using fused-silica for the mass. In particular, the lower coefficient of thermal expansion () leads to a small sensitivity to thermal gradients, which we see as a decrease in the (RX, RY) tilt drift. For a representative tilt resonance of , the tilt drift is for the fused-silica C-6D as opposed to for the aluminium precursor. The tilt drift of C-6D is limited by the epoxy bonding mechanism (see below) otherwise drift reduction would likely be even higher. Further improvements to the mass stability, particularly to temperaure fluctuations and thermal gradients, can be achieved by applying a gold coating to the silica mass to reduce its emissivity.
@@
We can place the centre of mass above the bending point (corresponding to negative in Eq. 1) and generate a gravitational anti-spring effect. This reduces the overall angular stiffness analogously to the inverted pendulum, which has a history of use within the GW community [21, 75]. In consequence, we can theoretically reach any resonant frequency regardless of the test mass properties by properly tuning the vertical CoM position. Micron-level fine-tuning can be achieved by adding/removing gram-scale balancing masses. In principle, this means that there is no penalty to tilt sensitivity of M-6D from the mechanical response, despite its smaller size compared to C-6D. The only reduction in sensitivity comes from the larger readout noise due to the latter’s factor of 2.5 longer lever arm. In practise, it is challenging to fine-tune the CoM offset for compact test masses. Due to M-6D’s significantly smaller moment of inertia, its tilt resonance is approximately 4 times more sensitive to the change in CoM height than C-6D, with only a difference separating a tilt frequency of from instability.
@@ IV Sensing and Actuation
The principal benefit of interferometric sensors is through their superior sensitivity compared to typical electronic readout schemes. An additional benefit of our sensors is in the robustness of the readout algorithm during initial alignment. The sensors have an operating range spanning several centimetres with, in principle, no cost to performance provided a laser with a sufficient modulation range can be found. The sensor’s linearity is invariant to angular fluctuations, with only a mild impact on sensitivity (inversely to fringe contrast, see Ref. [80]). Therefore, it is possible to align all the sensors in-air with relative ease. Our sensor is susceptible to strong laser frequency noise coupling. A balanced Michelson interferometer can strongly reject laser frequency noise. However, it is necessary for us to have a large arm-length imbalance to allow a strong coupling of the frequency modulation. However, as all of the sensing heads are connected to the same laser via an in-vacuum 8-way fibre splitter, the laser frequency noise is common to all sensors. Therefore, we place an extra reference sensing head inside the vacuum chamber providing a ‘null’ measurement of a fixed mirror. This gives us an independent readout channel for the laser frequency noise, which we use in feedback to the laser frequency modulation, and thus suppress this common noise source at the laser without impacting the individual sensing heads’ responses.
@@
We mitigate this issue through remote balancing. The presence of the DC force applied through the catching servo is a consequence of imperfect initial balancing and subsequent tilt drifts. The neutral position differs in-air due to buoyancy effects and so a significant one-off tilt shift can be introduced after pump-down with additional gradual drifts accumulating over time. As a result, this leads to an accumulation of DC compensating force. With in-vacuum remote balancing, we are able to shift the neutral balance towards the servo set point and thus manually zero out the DC force. We achieve this by shunting a small balancing mass with a series of magnetic pulses—a system we refer to as a magnetogun. The mass consists of a diameter, long brass rod with a neodymium magnet bonded on each end. The balancing rods are held on the test mass by friction and the magnetic pulses are injected on the millisecond-scale. This applies a large but short-duration force to overcome the static friction and delivers an impulse to the small balancing mass without exciting the dynamics of the comparatively heavy test mass. With this technique, we are able to move the masses in approximately increments, corresponding to a resolution of of DC force removal for C-6D. Under the tilt drift conditions specified in Sec. III, ignoring any sudden or seasonal changes to the environmental conditions, the rebalanced mass will not drift out of range of the DC actuators for over 10 years under normal conditions.
@@ V Sensitivity
We present the measurements of the ground motion together with the principal noise contributions in Fig. 2. Due to the similar parameters of the C-6D mass to the metal precursor, the sensitivity of most DoFs is consistent with those already presented in Ref. [15]. This is the result of a similar pendulum length, moments of inertia and interferometric sensor performance. In principle, our DFM-based sensors benefit from AC readout, meaning that certain noise sources (e.g. analogue-to-digital converter noise and laser intensity noise) are down-converted from a frequency region of lower relative noise and thus bypass the characteristic increase in the low-frequency noise (see discussion in Ref. [71]).
@@
The directly measured horizontal acceleration shown in Fig. 2(a) shows real motion larger than the readout noise floor across all frequencies. However, the readout nonlinearities do become dominant in the tilt-subtracted horizontal acceleration, limiting the sensitivity at . This noise can be substantially mitigated if the sensor is used in feedback within an inertial isolation scheme, as the nonlinear noise floor scales with the RMS readout signal (as shown in Ref. [79]). By reducing the RMS input motion, the nonlinearities decrease as well and the nominal sub-pm sensitivity of the interferometric sensors can be realised.
@@ Acknowledgements.
We acknowledge members of the LIGO Suspension Working Group for useful discussions, the support of the Institute for Gravitational Wave Astronomy at the University of Birmingham, STFC Equipment Call 2018 (Grant No. ST/S002154/1), STFC Consolidated Grant ”Astrophysics at the University of Birmingham” (No. ST/S000305/1), UKRI Quantum Technology for Fundamental Physics scheme (Grant No. ST/T006609/1 and ST/W006375/1), and UKRI ”The next-generation gravitational-wave observatory network” project (Grant No. ST/Y00423X/1). D.M. acknowledges the support of the 2021 Philip Leverhulme Prize.Appendix
@@
The tilt subtracted trace shows the result of a coherent subtraction of the horizontally-inferred tilt from the direct tilt measurement and is thus a good estimate of the residual noise below . For a better demonstration of how the tilt decoupling is performed in a full six-DoF inertial isolation of an active platform, see Ref. [14]. Above this frequency the subtracted signal becomes polluted by horizontal ground motion from the microseism. There is no tilt substracted trace for M-6D as the horizontally-inferred tilt is already below the sum of all noise sources. This is due, in part, to the worse sensitivity of M-6D but also due to the naturally larger ground tilt in the winter (C-6D measurement) compared to the spring (M-6D measurement).

### v2

@@ Abstract
We present a pair of seismometers capable of measurement in all six axes of rigid motion. The vacuum-compatible devices implement compact interferometric displacement sensors to surpass the sensitivity of typical electrical readout schemes. Together with the capability to subtract the sensitivity-limiting coupling of ground tilt into horizontal motion, our seismometers can widen the sensing band towards mHz frequencies. This has notable applications across a range of fields requiring access to low-frequency signals, such as seismology and climate research. We particularly highlight their potential application in gravitational-wave observatories (LIGO) in improving their observation capability of intermediate-mass black holes (). The sensors are based on a near-monolithic fused-silica design consisting of a fused-silica mass and fibre, showing improved stability and robustness to tilt drifts, alignment, and control compared to all-metal or mixed metal-silica designs. We demonstrate tilt sensitivity that surpasses the best commercial alternatives in a significantly reduced footprint compared to our previous iterations of these sensors.
@@ I Introduction
However, even more precise sensors are required for fundamental physics research, including observation of gravitational waves. In this case inertial sensors are used within a broader feedback scheme to sense and control the motion of platforms and larger devices—the technique of inertial isolation. Inertial isolation can also be seen as an essential enabling technology for quantum sensing and computing development, and applications in industrial manufacturing processes.
@@
In this paper, we present the design and performance of two six-axis inertial sensors based on ideas laid out in Ref. [12]. The purpose of these devices is to address the above limitations through (i) taking advantage of compact interferometric sensors with sub-picometer displacement sensitivity, (ii) creating an unconstrained system with full six-axis freedom of motion, and (iii) achieving low resonant frequencies along key DoFs. The ultimate goal is to provide better sensitivity across the full sensing band and to extend this band towards lower frequencies. The final designs and performance presented here follow extensive prior development of numerous metal precursors [13, 14, 15]. We enhanced the sensitivity and robustness (in terms of both mechanical and measurement stability) of the seismometers by (i) utilising fused silica instead of aluminium to suppress drifts, (ii) making the devices compact to fit into LIGO-type inertial isolation platforms, (iii) developing compact interferometers with deep-frequency modulation readout, and (iv) introducing ‘magnetoguns’ (see Section IV) to correct the angular position of the mass.
@@
These devices are versatile and thus can benefit a multitude of disparate fields of research. The sensor can be used directly as a seismometer, where it can detect low-frequency signals and thus obtain more complete information about seismic transients, as well as to provide insight into low-frequency seismic phenomena. The importance of low frequency seismic information has been demonstrated in earthquake analysis [16], hydrocarbon surveys [17], and climate research [18]. In all of these applications, the sensor would be effective due to its ability to decouple tilt from horizontal motion and its high sensitivity through the use of interferometric sensors. Better sensitivity can also reduce the amplitude of acoustic excitations generated during active seismic surveys, which would be less disruptive to local wildlife [19].
@@
Our inertial sensors can be integrated into actively controlled platforms and provide a vibration-free environment for general use. Often only a few DoFs are directly important for the given application. However, cross-couplings between the different DoFs inevitably lead to the most effective solution being to control all six DoFs regardless. Our sensor would be beneficial in such applications due to its high sensitivity and ability to measure all six DoFs in a single device. Most prominently for us is its use in gravitational wave (GW) detectors [20, 21, 22, 23] (see Section II). Additionally, our sensor can be applied to suppress vibrations of cryogenic systems [24, 25], electron microscopes [26] and used as a test-bed for space-borne instruments [27]. Outside of research applications, the sensors can be used in ultra-precision manufacturing [28] and photolithography [29, 30].
@@
Finally, the sensors we developed can be used in the role of a torsion balance. As part of our strict requirements for low-frequency sensitivity, we achieve low eigenfrequencies in the angular DoFs. This is achieved by adopting a design based on a torsion pendulum. Ultimately, we reach a level of sensitivity in the torsional DoF that is competitive with contemporary torsion balance experiments. Torsion balances have a strong history in gravity measurements with recent examples including Bosonic dark matter searches [31], and Newtonian noise measurement [32]. We have recently adapted the precursor 6D sensor [15] to investigate semi-classical gravity models [33, 34, 35], with results currently in preparation. It is also possible to apply such a device in air-borne gravity gradiometry [36, 37] and short-range force measurements investigating the Casimir effect in meta-materials [38, 39, 40].
@@ II LIGO Application
The Advanced LIGO [41] and Advanced Virgo [42] detectors have successfully yield dozens of GW sources [43, 44, 45] since the first GW transient detection in 2015 [46]. However, the duty cycle of the detectors, as well as their sensitivity below 25 Hz [47, 48], is significantly limited by ground vibrations (specifically, controls noise from the feedback loops used to control the position of the primary optics). The reduced duty cycle leads to the loss of otherwise detectable sources and the narrower detection band [49] limits the accumulation of signal-to-noise ratio. This degrades the parameter estimation and sky localisation, and limits the early warning time [50] for follow-up and multi-messenger observations [51]. Widening the band also increases the prospects to detect and study intermediate-mass black holes [52].
@@
Commercial sensors are limited by the noise of their inductive/capacitive readout exacerbated by the degrading response below . Horizontal single-axis sensors further suffer from tilt-horizontal cross-couplings. Ground tilt around the orthogonal horizontal axis couples into the readout with a factor of [60], where is the angular frequency, which masks the true horizontal signal at low frequencies. One solution is to make an independent estimate of the ground tilt with two spatially separated vertical sensors and subtract it from the horizontal sensor. This method is limited by the relatively short, metre-scale coherence length of the ground tilt and the space constraints on the typical isolated platform. It also relies on multiple stages of coherent subtraction, up to a common-mode rejection ratio of 1000, which is highly susceptible to calibration errors and systematics.
A wide range of custom inertial sensors have been demonstrated or proposed in the literature. Sensors with minimal tilt coupling have been proposed [61] to bypass the cross-coupling effect by, for example, investigating novel geometries and using multi-stage mechanical systems that offer negligible coupling to ground tilt. In the other direction, dedicated rotation sensors [62, 63, 64] have been developed, consisting of beams or cylinders with large moment of inertia, which offer high sensitive in tilt but negligible translational coupling. These sensors can be used to subtract the tilt signals that pollute conventional horizontal sensors. Improved sensors based around optical readout have also been proposed [65, 66, 67, 68, 69, 70]. These sensors are either variations upon the conventional single-axis design, or make sure of some novel geometry but, importantly, they demonstrate the strong application of optical (often interferometric) sensors in reaching substantial improvements in read-out sensitivity. Our focus has been on the development of a single six-axis seismometer that overcomes all of the key limitations simultaneously in a single device.
@@ III Mechanical Design
An essential component of our design is the single fused-silica suspension fibre—the same fibre in both tests. This fibre was pulled at the University of Glasgow similarly to the fibres on our precursor designs, using the same technique developed for the Advanced LIGO quadruple suspensions [72]. We observe the beneficial linear, elastic behaviour of the fibre, which exhibits no detectable hysteresis as witnessed in the RZ torsional motion. We witness this via the fibre returning to its initial position after responding to transient changes in conditions, such as pressure changes and thermal cycling during pump-down. In contrast, this property is notably different in other DoFs, such as tilt, where the mass shifts to a new mean position in response to such transients. This is due to the nonlinear behaviour of other materials, such as epoxy, which strongly influence those DoFs (see below). This, among other properties of fused silica, is a key motivation for our move towards a monolithic fused-silica design, replacing the previous aluminium test mass.
The fused-silica test mass is wheel-shaped with most of the mass concentrated at the edges of the ring to maximise the moment of inertia for the given mass. This allows us to reduce the actuation strength needed to control the linear DoFs and use a thinner fibre. The higher moment of inertia reduces the tilt resonances and thus leads to better tilt sensitivity. Here, we see the benefits of using fused-silica for the mass. In particular, the lower coefficient of thermal expansion () leads to a small sensitivity to thermal gradients, which we see as a decrease in the (RX, RY) tilt drift. For a representative tilt resonance of , the tilt drift is for the fused-silica C-6D as opposed to for the aluminium precursor. The tilt drift of C-6D is limited by the epoxy bonding mechanism (see below) otherwise drift reduction would likely be even higher. Further improvements to the mass stability, particularly to temperature fluctuations and thermal gradients, can be achieved by applying a gold coating to the silica mass to reduce its emissivity. This should not increase the thermal noise as the dominant thermal noise contribution originating from regions that store large elastic potential energy—the lossy epoxy layer connecting fibre to mass and the effective bending point of the fused silica fibre. Neither region will be gold-coated and hence will not increase the thermal noise.
@@
We can place the centre of mass above the bending point (corresponding to negative in Eq. 1) and generate a gravitational anti-spring effect. This reduces the overall angular stiffness analogously to the inverted pendulum, which has a history of use within the GW community [21, 75]. In consequence, we can theoretically reach any resonant frequency regardless of the test mass properties by properly tuning the vertical CoM position. Micron-level fine-tuning can be achieved by adding/removing gram-scale balancing masses. In principle, this means that there is no penalty to tilt sensitivity of M-6D from the mechanical response, despite its smaller size compared to C-6D. The only reduction in sensitivity comes from the larger readout noise due to the latter’s factor of 2.5 longer lever arm. In practise, it is challenging to fine-tune the CoM offset for compact test masses. Due to M-6D’s significantly smaller moment of inertia, its tilt resonance is approximately 4 times more sensitive to the change in CoM height than C-6D (i.e. derived from Eq. 1 is 4 times larger), with only a difference separating a tilt frequency of from instability.
@@ IV Sensing and Actuation
The principal benefit of interferometric sensors is through their superior sensitivity compared to typical electronic readout schemes. An additional benefit of our sensors is in the robustness of the readout algorithm during initial alignment. The sensors have an operating range spanning several centimetres. with, in principle, no impact on sensitivity. This is contingent upon choosing a laser with a sufficient modulation range to enable such large range. The sensor’s linearity is invariant to angular fluctuations, with only a mild impact on sensitivity (inversely to fringe contrast, see Ref. [80]). Therefore, it is possible to align all the sensors in-air with relative ease. Our sensor is susceptible to strong laser frequency noise coupling. A balanced Michelson interferometer can strongly reject laser frequency noise. However, it is necessary for us to have a large arm-length imbalance to allow a strong coupling of the frequency modulation. However, as all of the sensing heads are connected to the same laser via an in-vacuum 8-way fibre splitter, the laser frequency noise is common to all sensors. Therefore, we place an extra reference sensing head inside the vacuum chamber providing a ‘null’ measurement of a fixed mirror. This gives us an independent readout channel for the laser frequency noise, which we use in feedback to the laser frequency modulation, and thus suppress this common noise source at the laser without impacting the individual sensing heads’ responses.
@@
We mitigate this issue through remote balancing. The presence of the DC force applied through the catching servo is a consequence of imperfect initial balancing and subsequent tilt drifts. The neutral position differs in-air due to buoyancy effects and so a significant one-off tilt shift can be introduced after pump-down with additional gradual drifts accumulating over time. As a result, this leads to an accumulation of DC compensating force. With in-vacuum remote balancing, we are able to shift the neutral balance towards the servo set point and thus manually zero out the DC force. We achieve this by shunting a small balancing mass with a series of magnetic pulses—a system we refer to as a magnetogun. The mass consists of a diameter, long brass rod with a neodymium magnet bonded on each end. The balancing rods are held on the test mass by friction and the magnetic pulses are injected on the millisecond-scale. This applies a large but short-duration force to overcome the static friction and delivers an impulse to the small balancing mass without exciting the dynamics of the comparatively heavy test mass. With this technique, we are able to move the masses in approximately increments, corresponding to a resolution of of DC force removal for C-6D. Assuming the worst-case scenario that the steady-state tilt position will continue to drift unidirectionally by the observed over three months of operation indefinitely, the balanced mass will remain within range of the DC actuators for over 10 years. Given the desire to zero the DC actuation force, the mass can be periodically rebalanced with the magnetoguns. This may be at most once every few months to compensate for seasonal drifts in environmental conditions.
@@ V Sensitivity
We present the measurements of the ground motion together with the principal noise contributions in Fig. 2. Due to the similar parameters of the C-6D mass to the metal precursor, the sensitivity of most DoFs is consistent with those already presented in Ref. [15]. This is the result of a similar pendulum length, moments of inertia and interferometric sensor performance. Therefore, the reader should consult the results therein, in addition to the results highlighted below, for a comprehensible overview of the 6-axis performance.
In principle, our DFM-based sensors benefit from AC readout, meaning that certain noise sources (e.g. analogue-to-digital converter noise and laser intensity noise) are down-converted from a frequency region of lower relative noise and thus bypass the characteristic increase in the low-frequency noise (see discussion in Ref. [71]).
@@
The directly measured horizontal acceleration shown in Fig. 2(a) shows real motion larger than the readout noise floor across all frequencies. We obtain a tilt-subtracted curve by performing a coherent subtraction between the ‘true’ tilt DoF and inferred tilt from the horizontal DoF in the frequency region where the tilt-coupled motion dominates. This subtraction is straightforward owing to the simple coupling factor of between tilt and horizontal motion. We see a brief continuation of the true horizontal motion before reaching the sensitivity limit of the tilt DoF below . Here, the readout nonlinearities do become dominant, limiting the sensitivity at . These are the same nonlinearities, in origin and magnitude, to those we observed with the prototype design and can be most easily seen in the Z DoF as shown in Fig. 3(b) in Ref. [15]. The nonlinear noise floor is consistent with the predictions of the nonlinear noise model in fringe-counting interferometric sensors [79]. The nonlinear noise can be substantially mitigated if the sensor is used in feedback within an inertial isolation scheme, as the nonlinear noise floor scales with the RMS readout signal (as shown in Ref. [79]). By reducing the RMS input motion, the nonlinearities decrease as well and the nominal sub-pm sensitivity of the interferometric sensors can be realised.
@@ Acknowledgements.
We acknowledge members of the LIGO Suspension Working Group for useful discussions, the support of the Institute for Gravitational Wave Astronomy at the University of Birmingham, STFC Equipment Call 2018 (Grant No. ST/S002154/1), STFC Consolidated Grant “Astrophysics at the University of Birmingham” (No. ST/S000305/1), UKRI Quantum Technology for Fundamental Physics scheme (Grant No. ST/T006609/1 and ST/W006375/1), and UKRI “The next-generation gravitational-wave observatory network” project (Grant No. ST/Y00423X/1). D.M. acknowledges the support of the 2021 Philip Leverhulme Prize.Appendix
@@
The tilt subtracted trace shows the result of a coherent subtraction of the horizontally-inferred tilt from the direct tilt measurement and is thus a good estimate of the residual noise below . For a better demonstration of how the tilt decoupling is performed in a full six-DoF inertial isolation of an active platform, see Ref. [14]. Above this frequency the subtracted signal becomes polluted by horizontal ground motion from the microseism. There is no tilt subtracted trace for M-6D as the horizontally-inferred tilt is already below the sum of all noise sources. This is due, in part, to the worse sensitivity of M-6D but also due to the naturally larger ground tilt in the winter (C-6D measurement) compared to the spring (M-6D measurement).